OpenStudy (anonymous):

Solve for x and y. 5x-2y=8 -5x+6y=6 Help.

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

okay so you want to use process of elimination here giving you 4y=8

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

plug in y 5x-2(2)=8

OpenStudy (anonymous):

can you solve it now?

OpenStudy (anonymous):

Probably, I'll try it.

OpenStudy (anonymous):

uhhh... haven't you forgotten to add 8 + 6?

OpenStudy (anonymous):

yea I just realized that

OpenStudy (anonymous):

4y=14

OpenStudy (anonymous):

y= 7/2

OpenStudy (anonymous):

5x-2(7/2)=8

OpenStudy (anonymous):

my bad for forgetting that.

OpenStudy (anonymous):

So I got Y= 7/2 and X= 3, that's correct right?

OpenStudy (anonymous):

yup there you go good job

OpenStudy (anonymous):

It's ok. Thx for the help, peace.